### v1

# Towards Training-Free Open-World Classification with 3D Generative Models

Abstract
3D open-world classification is a challenging yet essential task in dynamic and unstructured real-world scenarios, requiring both open-category and open-pose recognition. To address these challenges, recent wisdom often takes sophisticated 2D pre-trained models to provide enriched and stable representations. However, these methods largely rely on how 3D objects can be projected into 2D space, which is unfortunately not well solved, and thus significantly limits their performance. Unlike these present efforts, in this paper we make a pioneering exploration of 3D generative models for 3D open-world classification. Drawing on abundant prior knowledge from 3D generative models, we additionally craft a rotation-invariant feature extractor. This innovative synergy endows our pipeline with the advantages of being training-free, open-category, and pose-invariant, thus well suited to 3D open-world classification.
Extensive experiments on benchmark datasets demonstrate the potential of generative models in 3D open-world classification, achieving state-of-the-art performance on ModelNet10 and McGill with 32.0% and 8.7% overall accuracy improvement, respectively. The code is available in the supplementary materials.
1 Introduction
Figure 1: Comparison of Pipelines: (a) SOTA methods project 3D samples into 2D images to harness 2D prior knowledge for open category classification but are sensitive to pose changes, (b) The proposed pipeline applies 3D prior knowledge to generate anchor samples for novel categories, embedding them in a rotation-invariant space for effective performance in open-pose scenarios.
3D classification within open-world environments is essential for developing computational models that can accurately interpret complex and unstructured real-world scenarios. Such advances are crucial for improving the autonomy and adaptability of intelligent systems Liu et al. (2023); Zhang et al. (2023); Zhao et al. (2024). This field aims to effectively recognize objects facing various real-world challenges, particularly those involving unexpected novel categories (open-category) and arbitrary object orientations (open-pose) Zhang et al. (2023); Zhu et al. (2023); Zhao et al. (2025).
Current efforts in open-world 3D classification rely primarily on transfer learning methods, categorized into language-only, language-point, and language-image model-based strategies Zhao et al. (2025). While these strategies showcase innovative methodologies, each encounters specific hurdles that inspire our investigation. Language-only model-based methods Cheraghian et al. (2020, 2022) rely solely on textual information to infer connections with novel classes, often fall short due to the limited descriptive power of text and the oversight of essential spatial information inherent in 3D data. Language-point model-based methods Xue et al. (2023, 2024); Xu et al. (2024), on the other hand, demands extensive data and significant computational resources for training, posing practical limitations. Language-image model-based methods Zhu et al. (2023); Huang et al. (2023); Hegde et al. (2023); Shen et al. (2024); Zhao et al. (2025) further complicate matters by requiring intricate projection hyperparameters to effectively map 3D data to 2D images. In particular, all existing approaches exhibit sensitivity to pose changes.
Recognizing these issues, we argue that the field currently faces a pressing need to effectively harness the spatial information inherent in 3D data while maintaining computational efficiency. In response to this challenge, we explore leveraging the rich prior knowledge of advanced 3D generative models Luo and Hu (2021); Nichol et al. (2022); Jun and Nichol (2023); Yi et al. (2024) by introducing an innovative pipeline that formulates a succinct and efficient 3D classifier for open-world tasks. This pipeline utilizes pre-trained 3D generative and representation models for object classification to achieve efficient performance without additional training, significantly reducing computational demands and data requirements. As shown in Fig. 1, the state-of-the-art (SOTA) methods employ language-image models, which project 3D samples into 2D images to exploit 2D prior knowledge for the classification of open categories. However, the generalizability of the 2D knowledge limits recognition, especially in open-pose scenarios where the object’s orientation and the projection style can affect the outcomes. Conversely, our proposed pipeline capitalizes on 3D prior knowledge to generate anchor samples for novel categories, subsequently embedding them into a rotation-invariant space to mitigate the 3D information loss inherent in projection settings.
The execution of our pipeline begins with feeding textual descriptions into a pre-trained text-to-3D generation model, which synthesizes anchor samples serving as prototypical examples for each category. Subsequently, a pre-trained point cloud representation model is utilized to extract anchor features from these samples. These anchor features are indicative of the classification capacity of our developed classifier, as its performance exhibits a positive correlation with the quality and diversity of these generated samples (see Section  4.3). Ideally, a sufficiently generalized generator will enhance the proficiency of our classifier in accurately categorizing any 3D objects. During the inference phase, cosine similarity within the representation space is computed to evaluate the similarity between test samples and anchor samples, thus leading to category predictions for the test samples. Further, we investigate the impact of various representation models on open-world classification. Our findings suggest that most representation models fail to provide consistent representations for identical 3D objects across different orientations, displaying rotational variance. Extensive experiments on the McGill and ModelNet10 datasets validate the robustness of our approach. Additionally, we explore the use of large language models (LLMs) to generate detailed textual descriptions of category names, with the aim of enhancing limited textual descriptive capabilities.
Our contributions can be summarized as follows:
  • We propose a training-free open-world pipeline for 3D classification. To our best knowledge, this is the first attempt to utilize a 3D generative model as prior knowledge in the open-world setting for such tasks.
  • We investigate the impact of the representation space within our pipeline on open-pose tasks and discuss the importance of rotation-invariant properties.
  • Utilizing 3D point cloud augmentation techniques, we have achieved state-of-the-art performance on ModelNet10 and McGill with overall accuracy improvement of 32.0% and 8.7%, respectively.
Figure 2: Overview of the network architecture with two components. (1) Open-world classifier: Category names are input into ChatGPT to create descriptions, utilized by a pre-trained text-to-3D model to generate anchor samples. These samples are augmented to align with test samples. A pre-trained 3D model extracts features, marking them as prototypes for classification. (2) Test sample inference: Test samples are augmented and processed through the same model for feature extraction. Predictions are based on cosine similarity between test and anchor sample features. Different colors denote new categories, gray represents test data.
2 Related Work
2.1 3D Open-World Classification
Current 3D open-world classification methods can be divided into three categories Zhao et al. (2025). Language-only model-based methods Cheraghian et al. (2020, 2022); Hao et al. (2023) utilize pre-trained language models like Word2Vec Mikolov et al. (2013) or GloVe Pennington et al. (2014) to establish semantic connections between training and novel classes but have fallen out of favor due to reliance on a single source of prior knowledge. Language-image model-based methods Zhu et al. (2023); Huang et al. (2023); Hegde et al. (2023); Shen et al. (2024); Zhao et al. (2025) project point clouds into multiple 2D views and utilize pre-trained models like CLIP Radford et al. (2021) or Diffusion Rombach et al. (2022); Sohl-Dickstein et al. (2015) for classifying these projections, enabling 3D open-world classification. Although effective, they involve numerous hyperparameters, such as projection angles and viewpoint weights, which complicate the process and diminish robustness. Language-point model-based methods Xue et al. (2023, 2024); Xu et al. (2024) link point clouds with text through various datasets and large language models. These methods need significant data resources and GPU memory, often with 7 or 13 billion parameter models. A recent study Zhao et al. (2025) highlights the limitations of existing methods in handling open-pose 3D objects in real-world applications. Conversely, our research initiates an innovative approach within the framework of language-point model methodologies by investigating current text-to-3D generative models Nichol et al. (2022); Jun and Nichol (2023); Yi et al. (2024) to address 3D open-world classification. Benefiting from the prior knowledge inherent in 3D generative models, our method achieves SOTA performance in open-pose scenarios without extensive hyperparameter tuning or large datasets for training.
2.2 Text-to-3D Generation
With the success of diffusion models in image generation Nichol et al. (2021), studies Luo and Hu (2021); Nichol et al. (2022); Jun and Nichol (2023); Yi et al. (2024) now integrate them into text-to-point-cloud tasks. Some methods Nichol et al. (2022); Jun and Nichol (2023); Yi et al. (2024) follow a ”text-to-image-to-point-cloud” paradigm, wherein a 2D image is first generated through the diffusion model and then converted into a 3D point cloud. Alternatively, some approaches Luo and Hu (2021); Ho et al. (2020) synthesize point clouds directly in 3D by sampling points and using a diffusion model. These pre-trained 3D generative models, rich in prior knowledge, can potentially serve as classifiers in open-world scenarios. We validate this using Shap-e Jun and Nichol (2023) and GaussianDreamer Yi et al. (2024), chosen for their speed and performance.
2.3 3D Represent Learning
Initial approaches to feature extraction from point clouds relied on geometric methods utilizing local information like normals and curvature Rister et al. (2017). PointNet Qi et al. (2017) was the first to use neural networks for spatial feature extraction. Inspired by the Transformer’s success in image and text domains Vaswani et al. (2017), many studies Wang (2023); Wu et al. (2024); Mehraban et al. (2024); Jiang et al. (2025) adopted its framework for 3D point cloud models, leveraging its attention mechanism to excel in capturing both local and global features. Addressing the open-pose challenge in open-world scenarios Zhao et al. (2025), we compare two models: TAP Wang et al. (2023), which lacks rotation invariance, and TET Melnyk et al. (2024), which tackles it. TET’s core idea involves a nonlinear TTlayer ($l_{TT}$), mapping point clouds from $\mathbb{R}^{N\times 3}$ to $\mathbb{R}^{N\times 4\times K}$ via steerable spherical neurons, embedding data into conformal space to ensure rotational and permutation equivariance.
3 Method
3.1 Problem Statement
Given $N_{c}$ categories with the names set $\mathbf{CN}=\{CN_{1},CN_{2},...,CN_{N_{c}}\}$, the 3D open-world classification task involves assigning the correct category label in $\mathbf{CN}$ to each instance within a sample set $\mathbf{TS}\in\mathbb{R}^{N_{t}\times N_{p}\times 3}$. Here, $N_{t}$ refers to the number of point cloud test samples, and $N_{p}$ implies the number of points in each sample. Typically, in an open-world setting, we are unable to obtain real samples corresponding to these categories to train a classifier.
3.2 Overall
The overall network architecture of the proposed method is depicted in Fig. 2. It consists primarily of two components: the construction of an open-world classifier and inferring test samples. Initially, we provide the given category text names as input to ChatGPT to retrieve comprehensive textual descriptions of each category. Subsequently, these textual descriptions are input into a pre-trained text-to-3D model to generate anchor samples corresponding to the respective categories. Data augmentation techniques are employed to process the anchor samples, thereby reducing the geometric representation differences between the anchor and test samples. Furthermore, we utilize the pre-trained 3D representation model to extract feature representations from these anchor samples. These extracted features serve as the prototypical examples for our open-world classifier. During the inference phase, we apply the same data augmentation procedures to the inference samples and input the augmented samples into the same representation model to obtain their feature representations. By comparing the cosine similarity between the feature representations of the test samples and the anchor samples, we can obtain category predictions for the test samples. More details follow in later subsections.
3.3 Construct the 3D Open-World Classifier
Usually, the category name $CN_{i}$ consists of only one or two words, they fail to offer sufficient guidance for generative models. Therefore, we propose employing LLMs to produce more detailed and informative descriptions $\mathbf{CD}$ of these category names. The procedure can be concisely summarized by the following formula:
$\mathbf{CD}=\mathbf{LLMs}\left(Q,\mathbf{CN}\right),$ (1)
where $Q$ is the prompt for LLMs, e.g.
What does the shape of an [$CN_{i}$] look like?” Specifically, when $CN_{i}$ is “Teddy Bear”, $Q_{i}$ would be “What does the shape of a teddy bear look like?”, and $CD_{i}$ would be “A teddy bear typically features a round head, large eyes, a small nose, and chubby body and limbs.”.
Anchor Sample Generation. The original or detailed textual descriptions, denoted as $\mathbf{CN}/\mathbf{CD}$, are input into text-to-3D generation models $\mathbf{GMs}$ to generate anchor samples $AS$ for each category. In the following text, all instances of $\mathbf{CD}$ can be replaced with $\mathbf{CN}$. In order to improve the robustness of the classifier we aim to construct, we generate $N_{a}$ anchor samples for each $CD_{i}$ by modulating the $noise_{j}$ within the pre-trained generative model. The process can be succinctly described by the following formula:
$AS_{i}^{j}=\mathbf{GMs}\left(noise_{j},CD_{i}\right),\quad AS_{i}^{j}\in\mathbb{R}^{N_{p}^{{}^{\prime}}\times 3},$ (2)
where $CD_{i}$ stands for the $i_{th},i\in(0,N_{c}]$ category text description, $AS_{i}^{j}$ serves as the $j_{th},j\in(0,N_{a}]$ anchor sample for the $i_{th}$ category, and $noise_{j}$ is the $j_{th}$ diffusion noise of text-to-3D generation models $\mathbf{GMs}$. In addition, $N_{p}^{{}^{\prime}}$ donates the number of points in the anchor sample.
Considering that there may be geometric representation differences between anchor samples and test samples, such as size and number of points included, we employ three standard data augmentation techniques: farthest point sampling (FPS), origin shifting, and scaling and rotation, to minimize these discrepancies. The process can be formally expressed as follows:
$\hat{AS_{i}^{j}}=\frac{\mathbf{FPS}(AS_{i}^{j},\hat{N_{p}})-\mathbf{Avg}(AS_{i}^{j})}{\mathbf{L}(AS_{i}^{j})}\cdot R,\quad R\in\mathbb{R}^{3\times 3},$ (3)
where $\mathbf{FPS}(\cdot,\hat{N_{p}})$ serves as the farthest point sampling operation, and $\hat{N_{p}}$ stands for the final number of sampling points. Additionally, $\mathbf{Avg}(\cdot)$ donates the operation to average coordinates of the point cloud, and $\mathbf{L}(\cdot)$ is to calculate the maximum Euclidean distance from any point in the point cloud to the centroid. Moreover, $R$ is the random rotation matrix.
Anchor Sample Feature Distribution. Different 3D representation models $\mathbf{RMs}$ have a significant impact on 3D open-world classifiers. Existing models are often limited to aligned datasets and overlook the influence of open-pose variations. After thorough exploration, we select TET Melnyk et al. (2024), a 3D representation model that focuses on the study of rotational variance, as the feature extractor for constructing our open-world classifier. To this end, the anchor sample feature $AF$ could be attained by the following formula:
$AF_{i}^{j}=\mathbf{RMs}\left(\hat{AS_{i}^{j}}\right),\quad AF_{i}^{j}\in\mathbb{R}^{\hat{N_{p}}\times D},$ (4)
where $AF_{i}^{j}$ serves as the $j_{th},j\in(0,N_{a}]$ anchor sample feature for the $i_{th}$ category, and $D$ donates the feature dimension. As shown in Fig. 2, the anchor sample feature distribution implies the classification ability of our 3D open-world classifier. The hyperplane separating the features acts as the decision boundary for our classifier. Furthermore, the region where the features of a test sample fall determines the predicted category of that test sample. In the following section, we offer a comprehensive explanation of how our constructed open-world classifier is employed to classify objects.
3.4 Inference Phase
Given the test sample set $\mathbf{TS}$, we apply the same data augmentation operations as those used for the anchor samples to minimize the discrepancy in geometric representations. The process can be succinctly described by the following formula:
$\hat{TS_{m}}=\frac{\mathbf{FPS}(TS_{m},\hat{N_{p}})-\mathbf{Avg}(TS_{m})}{\mathbf{L}(TS_{m})}\cdot R,\quad R\in\mathbb{R}^{3\times 3},$ (5)
where $\hat{TS_{m}}$ refers to the $m_{th},m\in(0,N_{t}]$ augmented sample. Accordingly, we input this feature into representation models $\mathbf{RMs}$ to obtain the corresponding test sample feature $TF_{m}$. The process can be formally expressed as follows:
$TF_{m}=\mathbf{RMs}\left(\hat{TS_{m}}\right),\quad TF_{m}\in\mathbb{R}^{\hat{N_{p}}\times D}.$ (6)
In order to measure the cosine similarity $d_{i}^{j}$ between the feature representation of the current test sample $TS_{m}$ and that of each anchor sample $AF_{i}^{j}$, we apply the following formula:
$d_{i}^{j}=1-\frac{AF_{i}^{j}\cdot TF_{m}}{\|AF_{i}^{j}\|_{2}\|TF_{m}\|_{2}},$ (7)
where $\|\cdot\|_{2}$ signifies the norm of the feature vector. Furthermore, the category name $CN_{k}$ associated with the anchor sample feature $AF_{k}^{j}$ that is most similar to the test sample feature $TF_{m}$ will be adopted as the predicted class $P_{m}$ for the test sample $TS_{m}$. This process is represented as follows:
$P_{m}=CN_{k},k=\arg\min_{i}(\{d_{1}^{1},d_{1}^{2},..d_{i}^{j}\}_{i,j=1,1}^{i,j=N_{c},N_{a}}),$ (8)
where $\arg\min_{i}(\cdot)$ denotes the index $i$ when the minimum value is taken.
4 Experiment
4.1 Experiment Setup
Datasets. The effectiveness of the proposed method is validated using the public datasets ModelNet40, ModelNet10, and McGill, which were introduced by OP3D Zhao et al. (2025).
ModelNet40/10 evolves from ModelNet40/10 Wu et al. (2015) through random rotation of the samples, resulting in open-pose variations. ModelNet40 is a benchmark dataset in the field of 3D point cloud classification. It is derived from computer-aided design models and contains high-quality 3D synthetic objects, free from noise and with regular structures. The dataset covers 40 categories of everyday objects (such as pianos, tables, airplanes, etc.) and consists of 12,311 3D objects, each represented by 10,000 points. The points are represented by 3D coordinates (x, y, z) and may include additional information such as normal vectors or colors (R, G, B). ModelNet10 is a curated subset of ModelNet40, focusing on 10 common object categories, such as beds, monitors, and chairs, containing a total of 4,899 3D objects.
McGill Zhao et al. (2025) is derived from McGill Siddiqi et al. (2008) by randomly rotating the samples in McGill and removing five categories that overlap with ModelNet40. McGill includes 14 categories of complex geometric shapes, such as ants, snakes, and hands. The objects in this dataset exhibit higher morphological diversity, particularly posing greater challenges when handling natural object categories. Each sample consists of point coordinates (x, y, z) and topological information.
Evaluation Metric. Following SOTAsZhang et al. (2023); Huang et al. (2023); Zhao et al. (2025), we evaluate the open-world classification performance using top-1 accuracy and mean accuracy (mAcc). Top-1 accuracy measures the proportion of samples for which the model’s top prediction matches the ground truth, while mAcc calculates the average accuracy across all categories to assess the balance of performance among classes.
Implement Details. All feature extraction and inference tasks were performed on an NVIDIA RTX 3090 GPU. For the point cloud diffusion model, we utilize a random seed $s\sim\mathcal{U}(0,50)$ to introduce diversity into the generated anchor point clouds of any category.
During the data augmentation stage, all the anchor samples and test samples are downsampled to 1024 points by the farthest point sampling. For the feature output dimensions of different models, the TET pre-trained model outputs features with dimensions of (256, 1), and the TAP pre-trained model outputs features with dimensions of (256, 384). To facilitate subsequent feature comparison and processing, we apply max pooling along the second dimension of the TAP model’s output, ultimately unifying the features to dimensions of (256, 1).
4.2 Performance Comparison
Our approach is compared against six SOTA methods: PointCLIP Zhang et al. (2023), ULIP Xue et al. (2023), ReconCLIP Qi et al. (2023), CLIP2Point Huang et al. (2023), PointCLIPv2 Zhu et al. (2023), and OP3D Zhao et al. (2025), on the open-pose and open-category classification datasets ModelNet10 and McGill. To ensure fairness, we utilized the pre-trained models from the official GitHub repositories provided by the respective papers (TAP and TET) and strictly followed the original experimental settings. The results presented here do not include the samples generated from LLM-generated prompts. That is, all anchor samples were directly generated using the category names. The impact of LLM-generated prompts on the results will be discussed in subsequent sections.
Our approach exhibits notable enhancements in classification accuracy over current models in both open-pose and open-category recognition tasks. The performance of our model on ModelNet10 dataset is shown in Table 1. Notably, we achieved a 32.0% improvement in overall accuracy (oAcc) and a 30.3% improvement in mean accuracy (mAcc). On McGill dataset, our model also demonstrated competitive performance. Compared to the current leading method, our approach improved overall accuracy (oAcc) by 8.7% and mean accuracy (mAcc) by 2.8%.
Table 1: Comparison to SOTAs. Proj. indicates whether it is necessary to project the 3D samples into 2D images. denotes that there is an overlap between the test categories and the training categories of the pre-trained model utilized.
Methods Proj. ModelNet10 McGill
[…]
PointCLIP Y 17.7 16.4 12.2 13.3
ULIP N 14.4 13.8 14.8 16.1
ReconCLIP Y 14.4 13.8 14.8 16.1
CLIP2Point Y 15.6 14.3 15.7 17.3
PointCLIPv2 Y 19.9 18.2 27.8 28.9
OP3D-CLIP Y 26.3 24.2 31.3 34.6
OP3D-Diffusion Y 22.6 21.7 39.1 44.7
Ours N 54.6 52.0 47.8 47.5
Figure 3: Comparison on the Number of Anchor Samples $N_{a}$.
Table 2: Comparison of classification effects of anchors generated by different generative models on ModelNet10.
[…]
Shap-e 22.0 19.0 66.0 31.4 10.5 65.0 67.4 79.0 65.0 87.0 53.5 51.2
GD 8.0 30.0 61.0 36.0 4.7 64.0 53.5 77.0 61.0 87.0 51.2 48.2
Ensemble 18.0 22.0 73.0 36.0 2.3 65.0 69.8 80.0 65.0 89.0 54.6 52.0
Table 3: Comparison of classification effects of anchors generated by different generative models on McGill. Abbreviations used: Din. = Dinosaur, Dol. = Dolphin, Oct. = Octopus, Qua. = Quadruple, Spe. = Spectacle, Spi. = Spider.
[…]
Shap-e 0.0 42.9 20.0 42.9 25.0 12.5 85.7 25.0 71.4 81.8 22.2 55.6 72.7 100.0 47.0 47.0
GD 20.0 28.6 0.0 57.1 25.0 25.0 14.3 12.5 71.4 90.9 22.2 33.3 45.5 100.0 39.1 39.0
Ensemble 0.0 42.9 20.0 57.1 25.0 25.0 57.1 25.0 71.4 90.9 33.3 44.4 72.7 100.0 47.8 47.5
Table 4: Comparison of classification effects of aligned-pose and open-pose settings on the different representation models using Shap-e.
[…]
ModelNet10 TAP 24.0 11.0 16.0 5.8 9.3 23.0 65.1 36.0 3.0 25.0 21.5 21.8
TET 22.0 19.0 66.0 31.4 10.5 65.0 67.4 79.0 65.0 87.0 53.5 51.2
ModelNet10 TAP 24.0 55.0 63.0 10.5 51.2 78.0 83.7 79.0 31.0 78.0 57.4 55.3
TET 24.0 16.0 69.0 29.1 33.7 69.0 41.8 79.0 67.0 84.0 53.5 51.3
4.3 Analysis of Results
The performance of the proposed method is influenced by several factors, including the number of anchor points, the selection of generative models, and 3D representation models. In this section, we will perform a quantitative analysis to evaluate the impact of each of these factors, thus demonstrating the effectiveness and robustness of our approach. The subsequent discussion will address each factor in detail.
Impact of Anchor Sample Size $N_{a}$.
We examined the impact of varying the number of anchors generated while fixing the dataset, represent model (TET), and generation model to evaluate its effect on classification performance. Three types of anchor generation approaches were tested: Shap-e, GaussianDreamer (GD), and their mixed (Ensemble).
As shown in Fig. 3, as the number of generated samples increased, the classification performance improved and stabilized. However, for the McGill dataset, performance slightly decreased at 8 anchor samples, probably due to individual sample differences causing fluctuations in the smaller anchor size. While more anchor samples generally improve performance, too many can lead to diminishing returns and increased generation time, necessitating a trad-off between performance and efficiency. Based on these findings, we set the number of anchor points to 7 for all the experiments.
Comparison on Generative Models In this experiment, we kept the dataset, represent model (TET), and the number of generated anchor points constant, while only changing the generative models to test their impact on classification performance. Two generative models were selected: Shap-e and GD. Both models can generate point clouds based on the given prompts, but they differ in methodology. Shap-e directly generates a point cloud, whereas GD first generates a coarse point cloud and then optimizes it using 3D Gaussian Splatting. In terms of generation time, Shape required 7 seconds, while GD took 15 minutes. The results of modelnet10 and mcgill are shown in Table 5.
The results on Modelnet10 are shown in Table 2. Overall, Shap-e outperforms GD in terms of classification accuracy, but it does not consistently lead in every category. Both generative models have their strengths: Shap-e excels in some categories, e.g. bathtub, chair, and monitor, while GD performs better in others, e.g. bed, desk, and table. Consequently, combining samples generated by both models results in improved performance. This shows that the diversity of samples generated by 3D generative models significantly affects the performance of the classification model.
Table 5: Performance comparison of classifiers using anchors generated by two 3D generative models, Shap-e and GaussianDreamer (GD), on ModelNet10 and McGill datasets.
Shap-e GD Accuracy
[…]
ModelNet10 53.5 51.2
51.2 48.2
54.6 52.0
McGill 47.0 47.0
39.1 39.0
47.8 47.5
Table 6: Comparison of prompts and the corresponding classification accuracies of samples generated using those prompts for ’dresser’.
Category Text Description $\mathbf{CD}$ oAcc(Dresser)$\uparrow$
A dresser. 10.5
A dresser with five drawers and circular handles. 4.7
An open design dresser with five exposed drawers. 7.0
A small bedside dresser with two drawers”. 12.8
A vintage dresser with carved details and two doors. 18.6
A wooden dresser with a built-in mirror and six drawers. 86.0
A decorative dresser with floral carvings and four drawers. 93.0
Table 7: Comparison of classification accuracy on the ModelNet10 using Shap-e with Category Name ($\mathbf{CN}$) and ChatGPT ($\mathbf{CD}$) prompts.
[…]
$\mathbf{CN}$ 22.0 19.0 66.0 31.4 10.5 65.0 67.4 79.0 65.0 87.0 53.5 51.2
$\mathbf{CD}$ 22.0 26.0 61.0 15.1 90.7 70.0 19.8 68.0 71.0 83.0 54.8 52.7
Comparison on Representation Models We compare two 3D representation models in this experiment: TAP, which lacks rotation invariance, and TET, which is designed to address this issue. The comparison is conducted on the aligned-pose dataset ModelNet10 and the open-pose dataset ModelNet10. In addition, we only utilize Shap-e as the generative model. Experimental results, shown in Table 4, demonstrate that on the aligned-pose dataset, TAP and TET exhibit comparable performance. Specifically, TET’s oAcc and mAcc are merely 0.1% higher than those of TAP. This implies that the two models, despite being designed with distinct approaches, display minimal performance disparities on the aligned-pose dataset. Notably, on the open-pose dataset, TET significantly outperforms TAP across the majority of categories. For instance, TET shows a significant improvement over TAP’s oAcc of 21.5% and mAcc of 21.8%.
It can be observed that designing specific structures to address point cloud rotation invariance effectively enhances the model’s robustness against rotation. The results show that TET, which incorporates such structures, is unaffected by point cloud rotations, as evidenced by the minimal difference in performance between the aligned-pose and open-pose datasets. In fact, TET performs slightly better on the open-pose dataset, with oAcc and mAcc increasing by 1.6% and 1.7%, respectively. In contrast, TAP, which uses a Transformer-based framework without specialized structures for rotation invariance, achieves results comparable to TET on the aligned-pose dataset. However, its performance drops sharply on the open-pose dataset, underscoring the need for rotation-invariant design to maintain robustness.
Figure 4: Visualization comparison of ”dresser” point clouds from original category names ($\mathbf{CN}$) and ChatGPT-generated prompts ($\mathbf{CD}$), which enhance anchors’ diversity.
4.4 Discussion on LLM-Generated Category Descriptions
During our experiments, we observed that the classification accuracy for certain categories remained consistently low. This was likely due to insufficient diversity in the generated samples. We hypothesized that generating more varied samples would improve the results. To address this, we explored optimizing descriptions for specific categories to improve classification performance. Taking the dresser category from the ModelNet10 dataset as an example, which had a notably low accuracy, we employed ChatGPT to generate multiple descriptive phrases tailored to the dresser category. These descriptions followed the format “a dresser that looks like … .” Using these optimized descriptions, we applied the Shap-e to create multiple new point clouds for the dresser category based on each description. The experimental results are presented in Tables 6, and the comparison between anchor points generated using ChatGPT-generated descriptions ($\mathbf{CD}$) versus category names ($\mathbf{CN}$) is shown in Fig. 4. The results demonstrate that appropriate descriptions significantly enhanced classification performance for the dresser category, while some inappropriate descriptions led to a decline in performance. However, since we did not know beforehand which descriptions would yield the best results, in subsequent experiments, we generated anchor samples using each description for all categories to further improve performance.
We applied this approach on the ModelNet10 dataset, generating LLM-generated category descriptions to create anchor samples for classification tasks. From the results, shown in Table 7, the oAcc and mAcc were increased by 1.3% and 1.5%, respectively. However, for many individual categories, we observed that the performance was worse compared to the results generated using only class names. We believe this is due to the higher variance in the samples generated by the model when using ChatGPT-generated descriptions, which covered a broader range of features. However, this also caused interference between categories, such as generating a dresser that closely resembled a table. Due to the instability of the descriptions produced by LLMs, we did not include these results in comparisons with other methods. Hence, improving control over LLM-generated descriptions for better results is a promising future direction.
5 Conclusion
In this paper, we propose a pipeline utilizes 3D point cloud generative models for 3D open-world classification. Unlinke existing methods that rely on large-scale 2D priors, our approach directly leverages the prior knowledge from 3D generative models. This completely avoids the 3D-to-2D projection process, thus mitigating the instability in open-pose scenes that could degrade performance due to projection settings. We also explore how the number and diversity of generated anchor samples, as well as rotation-invariant feature spaces, affect the proposed pipeline’s performance.
Experiments demonstrate SOTA performance on open-world datasets, highlighting the proposed pipeline as a promising avenue for future research in 3D open-world classification.